If the displacement, velocity and acceleration at an instant of a particle describing S.H.M are respectively 7.5m,7.5m/s,7.5m/s2. Calculate the maximum velocity of the particles


1
Expert's answer
2020-12-08T10:53:12-0500
"\omega^2=\frac{a}{x}=\frac{7.5}{7.5}=1\to \omega=1\frac{rad}{s}"

"V^2=v^2+\omega^2x^2=7.5^2+1^2(7.5)^2\\
V=10.6\frac{m}{s}"